OpenStudy (anonymous):

Precalc: Verify the identity. cos 4x + cos 2x = 2 - 2 sin2 2x - 2 sin2 x

OpenStudy (anonymous):

Is this your equation?\[\cos4x+\cos2x=2-2\sin^22x-2\sin^2x\]

OpenStudy (anonymous):

@SithsAndGiggles yes

OpenStudy (anonymous):

Okay. I think working with the left side would be easier. Use the following identity: \[\cos2x=\cos^2x-\sin^2x\] Note that the above identity also gives you \[\cos2x=1-2\sin^2x=2\cos^2x-1\] \[\cos4x+\cos2x=\cdots\\ (\cos^22x-\sin^22x)+(\cos^2x-\sin^2x)=\cdots\\ (1-2\sin^22x)+(1-2\sin^2x)=\cdots\]

OpenStudy (anonymous):

@SithsAndGiggles thanx :)

OpenStudy (anonymous):

You're welcome